I have 65 gallons of H2O, 35 gallons of 9% sulfuric acid.
How much of the sulfuric mix do I have to add to the H2O to get a 0.26% concentration

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

I have 65 gallons of H2O, 35 gallons of 9% sulfuric acid.
How much of the sulfuric mix do I have to add to the H2O to get a 0.26% concentration
--------------
Add S gallons.
65*0 + S*9 = (65 + S)*0.26
9S = 0.26S + 16.9
1.9336 gallons

Answer by stanbon(75887)

I have 65 gallons of H2O, 35 gallons of 9% sulfuric acid.
How much of the sulfuric mix do I have to add to the H2O to get a 0.26% concentration
--------------
Equation:
acid + acid = acid
0.09x + 0*65 = 0.0026(x+65)
----
Multiply thru by 10,000 to get:
900x = 26x + 26*65
-----
874x = 26*65
x = 1.934 gallons (amt of 9% to add)
-----------------------
Cheers,
Stan H.
===========